# Dipolar ordering transitions in many-body quantum optics: Analytical diagrammatic approach to equilibrium quantum spins

### v2

@@ I Introduction
One commonality of these quantum optical platforms is that their essential physics is often captured by many-body models whose microscopic degrees of freedom are quantum spins. However, the structure of the underlying spin-spin couplings is strikingly different to the short-range interactions encountered in solid-state based quantum magnetism [15]. This is due to long-wavelength photons and phonons that mediate the interactions between atoms and ions, which thereby give rise to tunable long-range and even all-to-all spin-spin couplings [16, 17, 18, 19, 20], see Fig. 1.
@@
Our goal is to investigate the properties of the equilibrium state of these models. We assume this state to be thermal and characterized in general by a finite temperature $T\geq 0$, see e.g. Ref. [21] for thermometry of a Rydberg-array [22]. This is certainly an approximation for quantum optical spin systems, which are well-isolated from the environment (in the sense that they are not embedded in a solid-state matrix). Nevertheless, when preparing ground states and also excited states there is a residual entropy, manifesting in fluctuations, albeit not necessarily thermal.
One recurrent quest in this field is to obtain phase diagrams which can be probed in experiment, e.g. in Rydberg tweezer arrays [22]. More recently, also dynamic quantities such as excitation spectra have come into reach [23]. Theoretically, these and other observables can be obtained from two-point spin correlators, see Eq. (7). For example, (second order) phase transitions are signaled by a divergence of the static correlator [24]. However, a large number of interaction partners per spin often causes considerable difficulties on the computational side: The plain high-temperature expansion of static correlators [25] becomes ineffective (see however Ref. [26]), and approaches relying on finite simulation volumes, e.g., density matrix renormalization group (DMRG) and quantum Monte Carlo (QMC), often suffer from finite-size effects. Furthermore, these methods may be negatively affected by high dimensionality ($d>1$) or the sign-problem [24].
@@
We build on foundations of the diagrammatic technique for quantum spins which were laid starting from the late 1960s [36, 37, 38, 39, 40, 41]. In Sec. II, we review the basic idea of the method and introduce a novel and efficient way to evaluate the diagrams to unprecedented order. Staying in the symmetry unbroken regime, in Sec. III we provide explicit expressions for the two-point functions of general $h=0$ Heisenberg models and the Ising and transverse-field Ising model (TFIM), the latter at $T=0$. In Sec. IV we benchmark our method on the hypercubic-lattice with nearest-neighbor interactions. Finally, in Sec. V, we apply our method to two particular quantum optical many-body systems with $S=1/2$, both on the square lattice: A power-law interacting ferromagnetic (FM) Heisenberg model at $T>0$ [42] and the Dicke-Ising model at $T=0$ [43, 44]. We conclude in Sec. VI.
@@ II.1 Models and perturbative expansion
with $\alpha_{1,2,3}\in\{x,y,z\}$ and the operator spin-length constraint $\mathbf{S}_{j}\cdot\mathbf{S}_{j}=S(S+1)$. The Hamiltonian is given in Eq. (1) where $S^{\pm}=(S^{x}\pm iS^{y})/\sqrt{2}$ are the spin ladder operators. In the following, it is understood that $\alpha\in\{x,y,z\}$ while $\gamma\in\{+,-,z\}$.
The following three choices for $h$ and $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}$ correspond to standard spin models that we consider in this work: (i) The Ising model (which is classical and treated mainly for reference)
@@
corresponds to $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}=\frac{1}{2}\left(1-\delta_{\gamma,z}\right)\left(1-\delta_{\gamma^{\prime},z}\right)J_{ii^{\prime}}$. Note that Ising models are usually defined for $S=1/2$. (iii) The Heisenberg model
@@
results from $h=0$ and $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}=\delta_{\bar{\gamma},\gamma^{\prime}}J_{ii^{\prime}}$ where we define $\bar{\gamma}$ by $\bar{+}=-$, $\bar{-}=+$ and $\bar{z}=z$. We keep $S$ general but assume vanishing onsite-coupling $J_{ii}=0$ for simplicity.
In the remainder of this work, we focus on the Matsubara spin-spin correlation function,
@@ II.2 Diagrams, J-reducibility and MF approximation
Next, the $V$-connected free spin average in Eq. (11) is taken by co-localizing the $2+2n$ spin operators (dots) in any possible way involving blocks of $m=2,3,4,...$ spins shown as brown ellipses, see Fig. 2(c) [41]. These blocks represent connected equal-site time ordered free spin correlators,
@@
which, for the homogeneous non-interacting Hamiltonian $H_{0,h}$ that we consider in Eq. (10), do not depend on the site index $i$. Connected spin correlators (with subscript $c$) are defined in analogy to their $V$-connected counterparts, see App. A. Due to our choice of $J_{ii}=0$, interaction lines cannot start and end at the same block. In Fig. 2(c), the absence of blocks of size $m=1$ representing single-spin averages (magnetization) is due to our current focus on the paramagnetic regime. The presence of blocks of order $m>2$ signals the absence of Wick’s theorem [27] for spin operators, ultimately rooted in the operator-valued right-hand side of the commutation relation (3) which differs from the canonical case of bosonic creation and annihilation operators.
The connected time ordered free spin correlators of Eq. (13), which are central diagrammatic building blocks will be discussed below in generality. For now, we mention the important special case that arises if all $m$ flavor indices are chosen as $z$ such that all involved operators $S_{i}^{z}(\tau_{n})=S_{i}^{z}$ commute. These connected z-spin correlators are then time-independent and given via the order-$(m-1)$ derivative of the Brillouin function $b_{c}(y)$ which describes the magnetization of a free spin $\left\langle S^{z}\right\rangle$ dependent on $y=\beta h$ [41],
@@
Given the above diagrammatic rules, it is straightforward to write down the 1JI diagrams for $\Sigma$ at order $n$ indicated by $\Sigma^{(n)}$. The diagrams for $n\leq 3$ which include the boxed diagrams of Fig. 2(b) are shown in Fig. 3. We label the different diagram topologies at order $n$ with an additional index $x=a,b,c,...$. Each diagram topology is uniquely identified by $(nx)$. The diagrams for $\Sigma^{(4)}$ have not been systematically collected so far in literature and are given in Fig. 7 of App. D.
Figure 3: Diagrams for $\Sigma$ up to order $J^{3}$ for a situation where single-spin expectation values are not involved. The blue multiplicity factors $\Lambda^{(nx)}$ denote the number of ways to arrive at a given diagram starting from the expansion in Fig. 2(b), e.g. in diagram (3b) there are three choices to pick the bottom line from the three lines available at third order perturbation theory. The two vertical lines are equivalent and do not enhance the multiplicity as an exchange can be compensated by a permutation of sites within the top brown ellipse.
We conclude this subsection by linking the diagrammatic expansion to the MF approximation for $G^{\alpha\alpha}$. The latter is obtained by using only the lowest-order approximation for $\Sigma$ (see the green arrow in Fig. 3)
@@
For $\alpha=z$, this evaluates to $b_{c}^{(1)}(\beta h)$, see Eq. (14). We also assumed $\left\langle S_{i}^{\alpha}\right\rangle=0$, otherwise $h$ in Eq. (17) would need to be replaced by the Weiss field, see Ref. [41] and Sec. V.2. The easiest way to see the correspondence between Eq. (17) and the MF approximation is to recall that the latter is exact for $N\rightarrow\infty$ if each spin interacts with all other spins in the system which requires to redefine $J\rightarrow J/N$. For $G$, this means that string-like diagrams as in the first line of Fig. 2(c) are the only finite contribution for $N\rightarrow\infty$. This follows since the site-indices of the internal (2-)blocks can be summed freely over all available sites only in string-like diagrams while being more restricted in diagrams that contain loops. Finally, the string-like diagrams are exactly the diagrams that are produced by the replacement $\Sigma\rightarrow\Sigma^{(0)}$ in the Larkin equation of Fig. 2(d). It follows that contributions $\Sigma^{(n)}$ with $n>0$ represent corrections to the MF results, see Figs. 1 and 3.
@@ II.3 Diagram evaluation via kernel functions
The remaining task is to calculate the analytical expressions encoded in the particular diagrams $\Sigma_{jj^{\prime}}^{\gamma\gamma^{\prime}(nx)}(i\nu_{m})$ which represent the different contributions of the right-hand side of Eq. (II.1). As the values for the blocks [brown ellipses, see Eq. (13)] are site-independent and we consider only a single type of spin-spin interaction $J$ in Eq. (1) (e.g. no XXZ interactions), we can combine all $J_{ii^{\prime}}$ [stemming from the insertions of $V$, c.f. Eq.(10)] along with the site sums in a geometry factor $t^{(nx)}_{jj^{\prime}}\left[J\right]\sim J^{n}$,
@@
To make progress, the crucial reformulation detailed in App. B expresses $\sigma^{(nx)}$ such that only the simple $\left\langle S^{\tilde{\gamma}_{1}}...S^{\tilde{\gamma}_{m}}\right\rangle_{0,c,h}$ are required. All the remaining complexity is encapsulated in the universal (model independent) kernel functions $K_{n+2}(\Omega_{1},...,\Omega_{n+2})$ of Ref. [48]. The latter originally appeared in the context of Fourier transforms of time-ordered correlation functions. The point is that an $n$-th order perturbative expression as in Eq. (II.1) can be understood as a Fourier transform of a time-ordered correlation function of order $n+2$ with $n$ frequencies set to zero.
@@
Expressions analogous to Eqns. (II.3) and (II.3) can be straightforwardly written for all diagram topologies and can be evaluated via computer algebra. In the next section, we provide the expressions for $\sigma^{\gamma\gamma^{\prime}(nx)}(i\nu_{m})$ obtained in this way. The sum over all topologies yields $\Sigma^{(n)}$ at order $J^{n}$,
@@ III Analytic results for 1-J irreducible Diagrams
In this section we report the analytic results of the diagram evaluation up to third order in $J$ for the Ising model, TFIM at $T=0$ and the Heisenberg model as defined in Eqns. (4), (5) and (6). Diagrams in fourth order and their analytic results are relegated to App. D. In the diagrams, single-spin blocks representing $\left\langle S_{i}^{\alpha}\right\rangle$ are not required. This is evident in the paramagnetic phase of the Ising and Heisenberg model where these quantities vanish. For the symmetry-unbroken phase of the TFIM they are finite for $\alpha=z$ but these blocks do not appear in the expansion of $G^{xx}=\mathrm{Re}[G^{++}+G^{+-}]$.
We provide results for the second contribution to Eq. (18), the geometry independent $\sigma^{\gamma\gamma^{\prime}(nx)}(i\nu_{m})$. For the Ising model and TFIM case we consider $\sigma^{zz}$ and $\sigma^{xx}$, respectively, see Tab. 2. In both cases diagrams which involve blocks of any odd order vanish and are not listed. For the Ising model this is due to spin flip symmetry $S_{i}^{z}\rightarrow-S_{i}^{z}\;\forall i$ in the paramagnetic phase, for the $\sigma^{xx}$ in the TFIM this follows from the fact that the blocks are solely made from $S^{+}$ and $S^{-}$ which need to appear in equal numbers for any finite contribution in light of U(1) spin rotation symmetry of $H_{0,h}$. For the Ising model only static contributions are finite due to its classical nature.
@@
Finally, we turn to the Heisenberg case at $T>0$ where the $\sigma^{zz(nx)}=\sigma^{\alpha\alpha(nx)}$ are reported in Tab. 3. We need to distinguish between the static case $\nu_{m}=0$ and the dynamic case $\nu_{m}=2\pi mT\neq 0$ for which we abbreviate $\Delta=\frac{1}{2\pi m}$. Again, these results have been tested for small spin clusters. Another non-trivial check for the resulting $\Sigma^{(n)}$ is the fulfillment of the property $\Sigma_{\mathbf{k}=0}^{(n)}(i\nu_{m}\neq 0)=0$ required by the constant-of-motion property of $S_{\mathbf{k}=0}^{z}$ in a Heisenberg system.
@@
We start with the Ising case, where $T_{c}$ was already obtained in Ref. [49] up to order $1/d^{3}$. We report the results for $T_{c}/T_{c}^{(0)}$ and $d=3,4,5,6,7$ in Tab. 5. For large $d$, the quasi-exact results from high-order series expansion or Monte Carlo simulations [51, 52, 53] are rapidly approached as the order $m$ in $1/d$ is increased. We also compare our results at order $1/d^{2}$ to other diagrammatic approaches with different resummation strategies: Our results are similar to the spin functional RG approach (spin-fRG) by Krieg and Kopietz [49], only at $d=3$ the latter has a slight advantage. The dynamical MF theory for spins [54] (spin-DMFT) is not competitive. This is not surprising given DMFT’s local approximation of $\Sigma$ already fails in order $J^{3}$ where the non-local diagram (3c) appears  [^1] Note that the lower-order non-local diagram (2b) vanishes for the Ising model.
@@
To showcase the advantage of analytical expressions for the Matsubara correlator, we consider the excitation gap $\Delta$ for $h\gtrsim h_{c}$ which we obtain from the dynamical $xx$-spin correlator, $G_{\mathbf{k}}^{xx}(i\nu\neq 0)$. The latter contains the dispersion $\omega_{\mathbf{k}}$ of spin waves transversal to the magnetization in the $S^{z}$ direction determined from the position of the sharp peak in $\mathrm{Im}G_{\mathbf{k}}^{xx,R}(\nu)$. For $h\rightarrow h_{c}$ from above, we expect $\omega_{\mathbf{k}}$ to vanish at $\mathbf{k}=\mathbf{N}$ where the spin wave softens and order sets in. For $h>h_{c}$, the gap is thus defined as $\Delta=\omega_{\mathbf{N}}$ and we consider
@@
$\Delta^{(1/d)}=\sqrt{{(h^{2}-2dShJ_{1}+\frac{5}{8}dSJ_{1}^{2})}/{(1+\frac{dSJ_{1}^{2}}{8h^{2}}})}.$ (35)
Setting $d=3$ and $S=1/2$, this agrees very well with recent iPEPS tensor network simulations in Ref. [59], see Fig. 4 (blue dashed line and dots). For $h\gtrsim h_{c}$, the iPEPS is actually closer to $\Delta^{(1/d)}$ than to $\Delta^{(1/d^{2})}$, we suspect this is an artifact of finite tensor-network bond dimension.
Figure 4: TFIM on the cubic lattice at $T=0$ and $S=1/2$. The vertical lines indicate the critical field $h_{c}$ in $1/d$-expansion Eq. (32), the quasi-exact result [58] (black) is rapidly approached with increasing expansion order. The same expansion is also used to estimate the spectral gap $\Delta$ shown by dashed lines. The first correction to MF approximation is already in good agreement to the tensor-network (iPEPS) simulation of Ref. [59].

### v3

@@ I Introduction
One commonality of these quantum optical platforms is that their essential physics is often captured by many-body models whose microscopic degrees of freedom are quantum spins. However, the structure of the underlying spin-spin couplings is strikingly different to the short-range interactions encountered in solid-state based quantum magnetism [15]. This is due to the spatially extended nature of the mode-functions of photons (phonons) that mediate the interactions between atoms (ions) and thereby give rise to tunable long-range and even all-to-all spin-spin couplings [16, 17, 18, 19, 20], see Fig. 1.
@@
Our goal is to investigate the properties of the equilibrium state of these models. We assume this state to be thermal and characterized by a finite temperature $T\geq 0$, see e.g. Ref. [21] for the demonstration of such a thermometry approach to recent experimental correlation data from a Rydberg-array experiment [22]. This is certainly an approximation for quantum optical spin systems, which are well-isolated from the environment (in the sense that they are not embedded in a solid-state matrix). Nevertheless, when preparing ground states and also excited states there is a residual entropy, manifesting in fluctuations, albeit not necessarily thermal.
One recurrent quest in this field is to obtain phase diagrams which can be probed in experiment, e.g. in Rydberg tweezer arrays [22]. More recently, also dynamic quantities such as excitation spectra have come into reach [23]. Theoretically, these and other observables can be obtained from two-point spin correlators, see Eq. (7) below. For example, (second order) phase transitions are signaled by a divergence of the static correlator [24].
On the computational side, however, a large number of interaction partners per spin often causes considerable difficulties: The plain high-temperature expansion of static correlators [25] becomes ineffective (see however Ref. [26]), and approaches relying on finite simulation volumes, e.g., density matrix renormalization group (DMRG) and quantum Monte Carlo (QMC), often suffer from finite-size effects. Furthermore, these methods may be negatively affected by high dimensionality ($d>1$) or the sign-problem [24].
@@
We build on foundations of the diagrammatic technique for general quantum spin-$S$ systems which were laid starting from the late 1960s [36, 37, 38, 39, 40, 41]. In Sec. II, we review the basic idea of the method and introduce a novel and efficient way to evaluate the diagrams to unprecedented order. For concreteness and simplicity, we specialize to specific types of spin-spin interactions in Eq. (1). In Sec. III we provide explicit expressions for the two-point functions of $h=0$ Heisenberg models and the Ising and transverse-field Ising model (TFIM), the latter at $T=0$. Throughout we stay in the symmetry unbroken regime for simplicity. In Sec. IV we benchmark our method on the hypercubic-lattice with nearest-neighbor interactions. Finally, in Sec. V, we apply our method to two particular quantum optical many-body systems with $S=1/2$, both on the square lattice: A power-law interacting ferromagnetic (FM) Heisenberg model at $T>0$ [42] and the Dicke-Ising model at $T=0$ [43, 44]. We conclude in Sec. VI.
@@ II.1 Models and perturbative expansion
with $\alpha_{1,2,3}\in\{x,y,z\}$ and the spin-length operator constraint $\mathbf{S}_{j}\cdot\mathbf{S}_{j}=S(S+1)$. The general bi-linear spin Hamiltonian with homogeneous magnetic field $h$ in z-direction is given in Eq. (1) where $S^{\pm}=(S^{x}\pm iS^{y})/\sqrt{2}$ are the spin ladder operators. In the following, it is understood that $\alpha\in\{x,y,z\}$ while $\gamma\in\{+,-,z\}$.
In the following, we specialize Eq. (1) to three important model classes defined by particular combinations of $h$ and the $3\times 3$ matrix $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}$ in flavor space. This specialization is not required by methodological restrictions but was chosen to keep the notation and diagrammatic complexity at a level suitable for presentation. Finally, in Sec. V.2 we consider a model with a more complicated structure.
The three simple choices correspond to the following standard spin models: (i) The Ising model (which is classical and treated mainly for reference)
@@
corresponds to $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}=\frac{1}{2}\left(1-\delta_{\gamma,z}\right)\left(1-\delta_{\gamma^{\prime},z}\right)J_{ii^{\prime}}$. Note that Ising models are usually defined only for $S=1/2$. (iii) The Heisenberg model
@@
results from $h=0$ and $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}=\delta_{\bar{\gamma},\gamma^{\prime}}J_{ii^{\prime}}$ where we define $\bar{\gamma}$ by $\bar{+}=-$, $\bar{-}=+$ and $\bar{z}=z$. We keep $S$ general and assume vanishing onsite-coupling $J_{ii}=0$ for simplicity.
In the remainder of this work, our computational focus will be on the Matsubara spin-spin correlation function,
@@
Given these considerations, in the following we will focus on this symmetry-unbroken phase and consider the Matsubara correlator (7) for $\alpha$ for which $\bigl{\langle}S_{i}^{\alpha}\bigr{\rangle}=0$. Among other simplifications on the diagrammatic side to be discussed below, this choice ensures the equality of connected and disconnected two-point correlators.
@@ II.2 Diagrams, J-reducibility and MF approximation
Next, the $V$-connected free spin average in Eq. (11) is taken by co-localizing the $2+2n$ spin operators (dots) in any possible way involving blocks of $m=2,3,4,...$ spins shown as brown ellipses, see Fig. 2(c) [41]. These blocks represent connected equal-site time-ordered free spin correlators,
@@
which, for the particular non-interacting Hamiltonian $H_{0,h}$ that we consider in Eq. (10), do not depend on the site index $i$. Connected spin correlators (with subscript $c$) are defined in analogy to their $V$-connected counterparts, see App. A. Due to our choice of $J_{ii}=0$, interaction lines cannot start and end at the same block. In Fig. 2(c), the absence of blocks of size $m=1$ representing single-spin averages (magnetization) is due to our current focus on the three models in question and on $G^{\alpha\alpha}$ for which $\bigl{\langle}S_{i}^{\alpha}\bigr{\rangle}=0$ in the symmetry-unbroken regime. The presence of blocks of order $m>2$ signals the absence of Wick’s theorem [27] for spin operators, ultimately rooted in the operator-valued right-hand side of the commutation relation (3) which differs from the canonical case of bosonic creation and annihilation operators.
The connected time-ordered free spin correlators of Eq. (13), which are central diagrammatic building blocks will be discussed below in generality. For now, we mention the important special case that arises if all $m$ flavor indices are chosen as $z$ such that all involved operators $S_{i}^{z}(\tau_{n})=S_{i}^{z}$ commute. These connected z-spin correlators are then time-independent and given via the order-$(m-1)$ derivative of the Brillouin function $b_{c}(y)$ which describes the magnetization of a free spin $\left\langle S^{z}\right\rangle$ dependent on $y=\beta h$ [41],
@@
Given the above diagrammatic rules, it is straightforward to write down the 1JI diagrams for $\Sigma$ at order $n$ indicated by $\Sigma^{(n)}$. The diagrams for $n\leq 3$ which include the boxed diagrams of Fig. 2(b) are shown in Fig. 3. We label the different diagram topologies at order $n$ with an additional index $x=a,b,c,...$. Each diagram topology is uniquely identified by $(nx)$. The diagrams for $\Sigma^{(4)}$ have not been systematically collected so far in literature and are given in Fig. 7 of App. D. Note that diagrams that combine the two external operators together with one internal operator in a $m=3$ block are 1JI by our definition. However, for the particular setup and observables that we focus on in this work (see above), these diagrams vanish.
Figure 3: Diagrams for $\Sigma$ up to order $J^{3}$ for the situation described in the text where magnetization is not relevant. The blue multiplicity factors $\Lambda^{(nx)}$ denote the number of ways to arrive at a given diagram starting from the expansion in Fig. 2(b), e.g. in diagram (3b) there are three choices to pick the bottom line from the three lines available at third order perturbation theory. The two vertical lines are equivalent and do not enhance the multiplicity as an exchange can be compensated by a permutation of sites within the top ellipse.
We conclude this subsection by linking the diagrammatic expansion to the MF approximation for $G^{\alpha\alpha}$. Quite simply, the latter is obtained by using only the lowest-order approximation for $\Sigma$ (green arrow in Fig. 3)
@@
For $\alpha=z$, this evaluates to $b_{c}^{(1)}(\beta h)$, see Eq. (14). We also assumed $\left\langle S_{i}^{\alpha}\right\rangle=0$, otherwise $h$ in Eq. (17) would need to be replaced by the Weiss field, see Ref. [41] and Sec. V.2. The easiest way to see the correspondence between Eq. (17) and the MF approximation is to recall that the latter is exact for $N\rightarrow\infty$ if each spin interacts with all other spins in the system which requires to redefine $J\rightarrow J/N$. For $G$ and in the limit $N\rightarrow\infty$, this means that string-like diagrams as in the first line of Fig. 2(c) are the only finite contribution. This follows since the site-indices of the internal ($m=2$-)blocks can be summed freely over all available sites only in string-like diagrams while being more restricted in diagrams that contain loops. Finally, the string-like diagrams are exactly the diagrams that are produced by the replacement $\Sigma\rightarrow\Sigma^{(0)}$ in the Larkin equation of Fig. 2(d). It follows that contributions $\Sigma^{(n)}$ with $n>0$ represent corrections to the MF results, see Figs. 1 and 3.
@@ II.3 Diagram evaluation via kernel functions
The remaining task is to calculate the analytical expressions encoded in the particular diagrams $\Sigma_{jj^{\prime}}^{\gamma\gamma^{\prime}(nx)}(i\nu_{m})$ which represent the different contributions of the right-hand side of Eq. (II.1). Two observations on the simple nature of the chosen models (Ising, TFIM and Heisenberg, c.f. Sec. II.1) are helpful but by no means crucial for progress: (a) The values for the blocks [brown ellipses, see Eq. (13)] are site-independent and (b) the coupling between any two blocks is characterized by a single number $J_{ii^{\prime}}$ (unlike for, say, XXZ interactions). Hence we can combine all $J_{ii^{\prime}}$ [stemming from the insertions of $V$, c.f. Eq.(10)] along with the site sums in a geometry factor $t^{(nx)}_{jj^{\prime}}\left[J\right]\sim J^{n}$,
@@
To make progress, the crucial reformulation detailed in App. B expresses $\sigma^{(nx)}$ such that only the simple $\left\langle S^{\tilde{\gamma}_{1}}...S^{\tilde{\gamma}_{m}}\right\rangle_{0,c,h}$ are required. All the remaining complexity is encapsulated in the universal (model independent) kernel functions $K_{n+2}(\Omega_{1},...,\Omega_{n+2})$ of Ref. [49]. The latter originally appeared in the context of Fourier transforms of time-ordered correlation functions. The key point is that an $n$-th order perturbative expression as in Eq. (II.1) can be understood as a Fourier transform of a time-ordered correlation function of order $n+2$ with $n$ frequencies set to zero.
@@
Expressions analogous to Eqns. (II.3) and (II.3) can be straightforwardly written for all diagram topologies and should be evaluated via computer algebra for efficiency. In the next section, we provide the expressions for $\sigma^{\gamma\gamma^{\prime}(nx)}(i\nu_{m})$ obtained in this way. Finally, the sum over all topologies yields $\Sigma^{(n)}$ at order $J^{n}$,
@@ III Analytic results for 1-J irreducible Diagrams
In this section we report the analytic results of the diagram evaluation up to third order in $J$ for the Ising model, TFIM at $T=0$ and the Heisenberg model as defined in Eqns. (4), (5) and (6). Diagrams in fourth order and their analytic results are relegated to App. D.
We provide results for the second contribution to Eq. (18), the geometry independent $\sigma^{\gamma\gamma^{\prime}(nx)}(i\nu_{m})$. For the Ising model and TFIM case we consider $\sigma^{zz}$ and $\sigma^{xx}$, respectively, see Tab. 2. In both cases diagrams which involve blocks of any odd order vanish and are not listed. For the Ising model this is due to spin flip symmetry $S_{i}^{z}\rightarrow-S_{i}^{z}\;\forall i$ in the paramagnetic phase, for the $\sigma^{xx}$ in the TFIM this follows from the fact that the blocks solely involve $S^{+}$ and $S^{-}$ which need to appear in equal numbers for any finite contribution in light of U(1) spin rotation symmetry of $H_{0,h}$. For the Ising model only static contributions are finite due to its classical nature.
@@
Finally, we turn to the Heisenberg case at $T>0$ where the $\sigma^{zz(nx)}=\sigma^{\alpha\alpha(nx)}$ are reported in Tab. 3. We need to distinguish between the static case $\nu_{m}=0$ and the dynamic case $\nu_{m}=2\pi mT\neq 0$ for which we abbreviate $\Delta=\frac{1}{2\pi m}$. Again, these results have been tested for small spin clusters. Another non-trivial check for the resulting $\Sigma^{(n)}$ is the fulfillment of $\Sigma_{\mathbf{k}=0}^{(n)}(i\nu_{m}\neq 0)=0$ required by the constant-of-motion property of $S_{\mathbf{k}=0}^{z}$ in a Heisenberg system.
@@
We start with the Ising case, where $T_{c}$ was already obtained in Ref. [50] up to order $1/d^{3}$. We report the results for $T_{c}/T_{c}^{(0)}$ and $d=3,...,7$ in Tab. 5. For large $d$, the quasi-exact results from high-order series expansion or Monte Carlo simulations [52, 53, 54] are rapidly approached as the order $m$ in $1/d$ is increased. We also compare our results at order $1/d^{2}$ to other diagrammatic approaches with different resummation strategies: Our results are similar to the spin functional RG approach (spin-fRG) by Krieg and Kopietz [50], only at $d=3$ the latter has a slight advantage. The dynamical MF theory for spins [55] (spin-DMFT) is not competitive. This is not surprising given DMFT’s local approximation of $\Sigma$ already fails in order $J^{3}$ where the non-local diagram (3c) appears  [^1] Note that the lower-order non-local diagram (2b) vanishes for the Ising model.
@@
To showcase the advantage of analytical expressions for the Matsubara correlator, we consider the excitation gap $\Delta(h)$ for $h\gtrsim h_{c}$ which we obtain from the dynamical $xx$-spin correlator, $G_{\mathbf{k}}^{xx}(i\nu\neq 0)$. The latter contains the dispersion $\omega_{\mathbf{k}}$ of spin waves transversal to the magnetization in the $S^{z}$ direction determined from the position of the sharp peak in $\mathrm{Im}G_{\mathbf{k}}^{xx,R}(\nu)$. The weight of the peak is the one-particle structure factor. For $h\rightarrow h_{c}$ from above, we expect $\omega_{\mathbf{k}}$ to vanish at $\mathbf{k}=\mathbf{N}$ where the spin wave softens and order sets in. For $h>h_{c}$, the gap is thus defined as $\Delta(h)=\omega_{\mathbf{N}}$ and we consider
@@
$\Delta^{(1/d)}(h)=\sqrt{{(h^{2}-2dShJ_{1}+\frac{5}{8}dSJ_{1}^{2})}/{(1+\frac{dSJ_{1}^{2}}{8h^{2}}})}.$ (35)
Setting $d=3$ and $S=1/2$, this agrees very well with recent iPEPS tensor network simulations in Ref. [60], see Fig. 4 (blue dashed line and dots). For $h\gtrsim h_{c}$, the iPEPS is actually closer to $\Delta^{(1/d)}(h)$ than to $\Delta^{(1/d^{2})}(h)$, we suspect this is an artifact of finite bond dimension of the tensor-network.
Figure 4: TFIM on the cubic lattice at $T=0$ and $S=1/2$. The vertical lines indicate the critical field $h_{c}$ in $1/d$-expansion Eq. (32), the quasi-exact result [59] (black) is rapidly approached with increasing expansion order. The same expansion is also used to estimate the spectral gap $\Delta(h)$ shown by dashed lines. The first correction to MF approximation is already in good agreement to the tensor-network (iPEPS) simulation of Ref. [60].